THE VISIBLE AND ULTRAVIOLET SPECTRUM OF SILVER CHLORIDE

\underline{LEAH O’BRIEN}, TENEIL KELLERMAN, and AMANDA LAMBETH, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The visible and ultraviolet electronic transitions of AgCl were recorded at high resolution with the Fourier transform spectrometer associated with the McMath-Pierce Solar Telescope at Kitt Peak, AZ. The ultraviolet system was previously observed and assigned as a 0^{+} - X ^{1}Σ^{+} transition,^{a} while the visible transition is a new, violet-degraded electronic system near 25,500 cm^{−1}. A similar, violet-degraded visible transition in AgF was recently observed,^{b} and high-level ab initio calculations^{c} on AgF have predicted that the excited state of this transition is a ^{3}Σ^{+}_{1} state. For the ultraviolet transition, the excited AgCl molecules were produced in a microwave discharge operated with 70 W absorbed power and at a total pressure of approximately 2 torr. For the visible system, a King-type carbon tube furnace was charged with AgCl powder and heated to 1900 C to produce the excited AgCl molecules. The molecular constants for the states involved will be presented.

aR.M. Clements and R.F. Barrow, Trans. Faraday Soc. 63, (1967) 2876-2878.

bH. Wang and J.L. Gole, J. Mol. Spectrosc. 161, (1993) 28-43.

cA. Ramirez-Solis and J. Schamps, J. Chem. Phys. 102, (1995) 4482-4490.